# On regions of mixed unitarity for semigroups of unital quantum channels

### v1

@@ 5. Behavior of dynamical semigroup of quantum channels near origin
$=\frac{1}{2}\big(\left\lvert\langle(I_{3}\otimes B^{2})\Omega_{3},\Omega_{3}\rangle\right\rvert^{2}+\frac{1}{9}\langle\Omega_{3},\Omega_{3}\rangle\big)=\frac{1}{2}\big(\frac{1}{9}\left\lvert\operatorname{\textsf{tr}}(\overline{B}^{2})\right\rvert^{2}+\frac{1}{9}\big)=\frac{1}{9}$
@@
$=(1-e^{-\frac{3}{2}t})\frac{1}{9}+e^{-t}(\frac{10}{18}b_{t}+\frac{1}{9}d_{t})$
@@
$=\frac{e^{-t}}{9}\big\{e^{t}-e^{-\frac{t}{2}}-\frac{4}{2}(\sinh(\frac{t}{2})-\sin(\frac{t}{2}))-\sinh(\frac{t}{2})\big\}\qquad(\because b_{t}=-\sinh(\frac{t}{2})-\sin(\frac{t}{2}))$
@@
6. Semigroups of mixed Weyl-unitary channels
A mixed unitary channel is a convex combination of maps of the form $\mathrm{Ad}_{U}$, where $U$ can be arbitrary unitary. Let $\mathscr{S}$ be a subset of the group of all $d\times d$ unitaries $\mathbb{U}(d)$. One can ask, What are all the channels that can be written as convex combinations of maps $\mathrm{Ad}_{U}$, where $U\in\mathscr{S}$? In this context we recall the notion of Weyl unitaries.
@@ 6. Semigroups of mixed Weyl-unitary channels
The operators $W_{a,b}$ are known as Weyl unitaries. The usefulness of Weyl unitaries in quantum information theory is well-known (See [Wat18]). They have the following key properties:
@@
The orthogonality condition tells us that, the set $\{\frac{1}{\sqrt{d}}W_{a,b}:a,b\in\mathbb{Z}_{d}\}$ is an orthonormal basis for $\mathbb{M}_{d}$ w.r.t. the Hilbert-Schmidt inner product. For any $a,b,a^{\prime},b^{\prime}\in\mathbb{Z}_{d}$, define $\phi_{(a,b),(a^{\prime},b^{\prime})}$ on $\mathbb{M}_{d}$ as
@@
[Wat18] Let $\Phi$ be a unital quantum channel on $\mathbb{M}_{d}$. Then the following are equivalent:
@@
$(ii)$ Since $\{\frac{1}{\sqrt{d}}W_{a,b}:a,b\in\mathbb{Z}_{d}\}$ is an orthonormal basis of $\mathbb{M}_{d}$ w.r.t. the Hilbert-Schmidt inner product, we have
@@
$(ii)$ Proof follows similar to the proof of 2.7 (iii). ∎
@@
$(iii)\Rightarrow(i)$ Assume that $\mathcal{L}$ is a Weyl-covariant linear map. Observe that if $\mathcal{L}$ is Weyl-covariant, then $\mathcal{L}^{k}$ is also Weyl-covariant. Therefore, for every $a,b\in\mathbb{Z}_{d}$ and $X\in\mathbb{M}_{d}$ we have
$e^{t\mathcal{L}}(W_{a,b}XW_{a,b}^{*})$ $=\lim_{n\to\infty}\sum_{k=1}^{n}\frac{t^{k}\mathcal{L}^{k}(W_{a,b}XW_{a,b}^{*})}{k!}$
$=\lim_{n\to\infty}\sum_{k=1}^{n}W_{a,b}\frac{t^{k}\mathcal{L}^{k}X)}{k!}W_{a,b}^{*}$
$=W_{a,b}\big(\lim_{n\to\infty}\sum_{k=1}^{n}\frac{t^{k}\mathcal{L}^{k}X)}{k!}\big)W_{a,b}^{*}$
$=W_{a,b}e^{t\mathcal{L}}(X)W_{a,b}^{*}.$
Thus $e^{t\mathcal{L}}$ is Weyl-covairant for every $t\geq 0$ and by Proposition 6.2 it follows $e^{t\mathcal{L}}$ is mixed Weyl-unitary for every $t\geq 0$. ∎
@@
Proof is similar to the Lemma 2.7. ∎

### v2

@@ 5. Behavior of dynamical semigroup of quantum channels near origin
$=\frac{1}{2}\big(\left\lvert\langle(I_{3}\otimes B^{2})\Omega_{3},\Omega_{3}\rangle\right\rvert^{2}+\frac{1}{9}\langle\Omega_{3},\Omega_{3}\rangle\big)=\frac{1}{2}\big(\frac{1}{9}\left\lvert\operatorname{\textsf{tr}}(\overline{B}^{2})\right\rvert^{2}+\frac{1}{9}\big)=\frac{1}{9}.$
@@
$=(1-e^{-\frac{3}{2}t})\frac{1}{9}+e^{-t}(\frac{10}{18}b_{t}+\frac{1}{9}d_{t})\qquad(\because\langle\operatorname{\textsf{id}},\Gamma\rangle=0)$
@@
$=\frac{e^{-t}}{9}\big\{e^{t}-e^{-\frac{t}{2}}-\frac{4}{2}(\sinh(\frac{t}{2})+\sin(\frac{t}{2}))-\sinh(\frac{t}{2})\big\}\qquad(\because 2b_{t}=-\sinh(\frac{t}{2})-\sin(\frac{t}{2}))$
@@
6. Semigroups of mixed Weyl-unitary channels and Schur channels
In this Section we study some special classes of channels which are of interest in quantum infomration theory. A mixed unitary channel is a convex combination of maps of the form $\mathrm{Ad}_{U}$, where $U$ can be arbitrary unitary. Let $\mathscr{S}$ be a subset of the group of all $d\times d$ unitaries $\mathbb{U}(d)$. One can ask, What are all the channels that can be written as convex combinations of maps $\mathrm{Ad}_{U}$, where $U\in\mathscr{S}$? In this context we recall the notion of Weyl unitaries.
@@ 6. Semigroups of mixed Weyl-unitary channels and Schur channels
The operators $W_{a,b}$ are known as Weyl unitaries. The usefulness of Weyl unitaries in quantum information theory is well-known (See [39]). They have the following key properties:
@@
The orthogonality condition tells us that, the set $\{\frac{1}{\sqrt{d}}W_{a,b}:a,b\in\mathbb{Z}_{d}\}$ is an orthonormal basis for $\mathbb{M}_{d}$ with respect to the Hilbert-Schmidt inner product. For any $a,b,a^{\prime},b^{\prime}\in\mathbb{Z}_{d}$, define $\phi_{(a,b),(a^{\prime},b^{\prime})}$ on $\mathbb{M}_{d}$ as
@@
[39] Let $\Phi$ be a unital quantum channel on $\mathbb{M}_{d}$. Then the following are equivalent:
@@
$(ii)$ Since $\{\frac{1}{\sqrt{d}}W_{a,b}:a,b\in\mathbb{Z}_{d}\}$ is an orthonormal basis of $\mathbb{M}_{d}$ with respect to the Hilbert-Schmidt inner product, we have
@@
$(ii)$ Proof follows similar to the proof of 2.6 (iii). ∎
@@
$(iii)\Rightarrow(i)$ Assume that $\mathcal{L}$ is a Weyl-covariant linear map. This implies that $\mathcal{L}^{k}$ is Weyl-covariant for every $k\geq 0$. Then by power series expansion, $e^{t\mathcal{L}}$ is Weyl-covairant for every $t\geq 0$ and by Proposition 6.2 it follows $e^{t\mathcal{L}}$ is mixed Weyl-unitary for every $t\geq 0$. ∎
@@
The proof is similar to that of Lemma 2.6. ∎
@@
Now we consider another important class of channels called Schur channels. Recall that a CP map $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$ is said to be a Schur channel if there exists a $C\in\mathbb{M}_{d}^{+}$ with unit diagonal entries such that
$\Phi(X)=\mathbb{S}_{C}(X):=C\circ X,$
where $C\circ X$ is the Hadamard (or Schur) product of $C$ and $X$. A positive matrix $C\in\mathbb{M}_{d}^{+}$ with unit diagonal entries is referred as a correlation matrix. It is straightforward to verify that, a CP map $\Phi$ is a Schur channel if and only if it admits a Kraus representation with all Kraus operators being diagonal. Since the linear span of Kraus operators of a CP map is independent of the representation it follows that in any Choi-Kraus representation of a Schur channel all Kraus operators are diagonal.
Definition 6.9.
A correlation matrix is said to be a mixed rank-one matrix if it is a convex combination of rank one correlation matrices.
Observe that a rank one correlation matrix is of the form $B=[e^{i(b_{j}-b_{k})}]_{1\leq j,k\leq d}$ for some real numbers $b_{1},b_{2},\ldots,b_{d}.$ In such a case, $\mathbb{S}_{B}=\mathrm{Ad}_{U}$, where $U$ is the diagonal matrix with entries $e^{-ib_{1}},e^{-ib_{2}},\ldots,e^{-ib_{d}}.$ Moreover, it is well-known that a Schur channel $\mathbb{S}_{C}$ is mixed unitary if and only if $C$ is a mixed rank-one correlation matrix. (See the proof of Lemma 2.4 in [8] or Theorem 5 [30]). Schur channels which are not of this form known to exist for $d\geq 4.$
The correlation matrices are of interest in quantum theory due to their connection with Bell’s inequality and related concepts ([5]). They are of interest in quantum information theory as they provide very simple and natural examples of unital quantum channnels ([39]). There exists extensive literature on correlation matrices and associated quantum channels. Nevertheless the following result on eventual mixed rank-one property seems to be new.
Theorem 6.10.
Let $C=[c_{jk}]_{1\leq j,k\leq d}$ be a correlation matrix. Then $\{[c_{jk}^{n}]:n\in{\mathbb{N}}\}$ is eventually mixed rank-one, that is, there exists $K\in\mathbb{N}$ such that $c_{jk}^{n}]$ is mixed rank-one for all $n\geq K.$ Similarly, if $A=[a_{jk}]_{1\leq j,k\leq d}$ is a matrix such that $e^{ta_{jk}}]:t\geq 0\}$ are correlation matrices, then there exists $T\geq 0$ such that $e^{ta_{jk}}]$ is mixed rank-one for all $t\geq T.$
Proof.
The first part is an immediate consequence of our main Theorem 3.9, by considering the CP channel $\mathbb{S}_{C}$ along with the observation made above that if a Schur channel of a matrix is mixed unitary then the matrix is a mixed rank-one matrix. The second part follows from the one parameter semigroup version of the main theorem proved in Theorem 4.12. ∎
Now we ask when does a semigroup of Schur channels remain mixed unitary for all times? Since a CP map is a Schur channel if and only if it admits a Kraus representation consisting of diagonal operators, this question is equivalent to determining when a semigroup of unital channels is $G$-mixed unitary for all times, where $G$ denotes the group of diagonal unitary matrices.
Theorem 6.11.
Let $A=[a_{jk}]\in\mathbb{M}_{d}.$ Consider the semigroup of linear maps $\Phi_{t}=\mathbb{S}_{C(t)}$, where $C(t)=[e^{ta_{jk}}]\in\mathbb{M}_{d}$. Then the following are equivalent:
  1. (1)
    $\Phi_{t}$ is mixed unitary for all $t\geq 0$;
  2. (2)
    $C(t)$ is mixed rank-one for all $t\geq 0$;
  3. (3)
    The matrix $A$ is in the closed cone generated by $\{[e^{i(b_{j}-b_{k})}-1]_{1\leq j,k\leq d}:b_{j}\in\mathbb{R},\penalty 10000\ \forall j\}$;
  4. (4)
    AA is a positive linear combination of matrices of the following form: (i) $i(h_{j}-h_{k})]_{1\leq j,k\leq d}:h_{j}\in\mathbb{R},\penalty 10000\ \forall j\}$; (ii) $\{[-(a_{j}-a_{k})^{2}]_{1\leq j,k\leq d}:a_{j}\in\mathbb{R},\penalty 10000\ \forall j\};$ (iii) $\{[e^{i(b_{j}-b_{k})}-1]_{1\leq j,k\leq d}:b_{j}\in\mathbb{R},\penalty 10000\ \forall j\}$.
Proof.
We have already seen the equivalence of $(1)$ and $(2).$ Let $\Phi_{t}=\mathbb{S}_{C(t)}$ be the given semigroup of linear maps. Then its generator $\mathcal{L}$ is given by $\mathcal{L}=\mathbb{S}_{A}$. Let $G$ be the group of diagonal unitary matrices. Observe that, for any $t\geq 0$, $\Phi_{t}$ is mixed unitary if and only if $\Phi_{t}$ is $G$-mixed unitary. Therefore, from Theorem 6.8, $\Phi_{t}$ is mixed unitary for all times if and only if $\mathcal{L}$ is in the closed convex cone generated by $\{\mathrm{Ad}_{U}-\operatorname{\textsf{id}}:U\in G\}$. Since $U$ is diagonal the map $\mathrm{Ad}_{U}$ is a Schur channel of the form $\mathbb{S}_{[e^{i(b_{j}-b_{k})}]}$. As the map $\mathbb{M}_{d}\ni B\mapsto\mathbb{S}_{B}$ is linear, $(3)$ follows. Now from [25] or Theorem 4.3, the generator of the semigroup has the form
$\mathbb{S}_{C}(X)=i[H,X]+\frac{1}{2}\sum_{j=1}^{n}\lambda_{j}(2L_{j}^{*}XL_{j}-L_{j}^{*}L_{j}X-XL_{j}^{*}L_{j}),$
with self-adjoint $H$ and either self-adjoint or unitary $L_{j}^{\prime}s$ and $\lambda_{j}\geq 0.$ Since as Schur property forces $H$, $L_{j}$’s to be diagonal we get $(4)$. The implication $(4)$ implies $(1)$ is clear from Theorem 4.3. ∎
@@ Acknowledgements
Data availability
The manuscript has no associated data.
Declarations
Conflict of interest
The authors have no conflict of interest.